Fermi National Laboratory


Accelerator Update for August 1, 2005 - August 3, 2005

Monday August 1
The day shift began with Operations monitoring store 4312 and a Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, with MiniBooNE, NuMI, taking beam, and with no beam to the Meson area due to beam line work.

Booster experts began work on a kicker (MKS03) at 8:26 AM; they completed the work about ninety minutes later.

A Main Injector (MI) expert began a slipstacking study at 9:46 AM.

Operations terminated store 4312 at 10:34 AM. Operations then put the Tevatron (TeV) into shot setup.

Operations established combination store 4313 at 1:32 PM with an initial luminosity of 104.27E30.

Pbar resumed stacking at 1:48 PM.

At 4:41 PM, NuMI tripped off for the first of seven times in about three hours. It reset each time.

Operations put the Recycler into shot setup around 5:23 PM to stash antiprotons. Operators completed the stash at 6:23 PM.

Pbar resumed stacking at 6:38 PM.

Operations established beam to the Meson experiments MCenter and MTest at 8:51 PM.

MTest experiments reported at 11:55 PM that they were done with the beam for the night.

Tuesday August 2
The midnight shift began with Operations monitoring store 4313 and the Recycler stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Recycler began stashing antiprotons at 12:31 AM. The stashing ended and Pbar resumed stacking around 1:53 AM.

MCenter resumed taking beam at 1:55 AM.

At 8:03 AM, Operations received a TeV sector C1 air pressure low alarm. Operators responded and fortunately got the air compressor back on before the air pressure dropped enough to cause a beam line trip, which would have aborted the store.

At 10:21 AM, the antiproton production and stacking rates went to zero. Operations and experts investigated.

At 11:23 AM, a NuMI power supply (HV101) tripped on a ground fault indication and took NuMI down. It wouldn’t reset. An expert investigated and reported that he had the supply back by 12:09 PM, but that he couldn’t find a problem with the supply. NuMI resumed taking beam at 12:17 PM.

Operations reported at 1:22 PM that Pbar experts managed to resume near-normal stacking rates. The expert then began an optic study for beam traveling from P1 to AP1.

Pbar experts reported at 1:41 PM that the stacking problem was due to an AP-50 vacuum problem.

Operations terminated store 4313 at 7:07 PM.

The Recycler electron cooling people reported a controls problem at MI-31. Experts turned off the Pelletron while they investigated. Operations attempted to use stochastic cooling on the antiprotons, but the beam still looked bad longitudinally.

Operations put the TeV through a dry squeeze at 7:38 PM.

Operations put the TeV into shot setup at 9:25 PM.

Operations established combination store 4315 at 10:42 PM with an initial luminosity of 102.8E30. Pbar resumed stacking.

MCenter resumed taking beam at 10:57 PM.

Wednesday August 3
The midnight shift began with Operations monitoring store 4315 and a small stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

The Recycler began stashing antiprotons at 2:31 AM. The stash ended and Pbar resumed stacking at 3:37 AM.

MCenter resumed taking beam at 3:52 AM.

Operations put the Recycler into shot setup at 8:03 AM.

The Plans for Wednesday
The plan for today is to stack, stash, and store. MI experts will be allowed to conduct a beam matching study.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/3/2005   email Fermilab